OpenStudy (anonymous):

A(t)=-16t^2+120t what is the maximum height of the arrow

OpenStudy (anonymous):

A(t)=-16t^2+120t = +120t - (4t)^2 = 225 -225 +2(4t)*15 -(4t)^2 = 225 - (15-4t)^2

OpenStudy (anonymous):

hope u can do from here ... max value of A(t) is 225

OpenStudy (loser66):

I think, that is the downward parabola, so that the maximum is the vertex of it. And we have formula to find out the vertex. It is x-coordinate = \(\dfrac{-b}{2a}\) then plug back to the original equation to find A that is the y-coordinate of the vertex.